Part of the proceeds from a garage sale was $460 worth of $10 and $20 bills. If there are 10 more $10 bills than $20 bills, find the number of each denomination.
Answer by addingup(3677)

x= y+10
10x+20y= 460
10(y+10)+20y= 460
10y+100+20y= 460
30y= 360
y= 12 You have 12 $20 bills and the problem says you have 10 more $10 bills than $20 bills:
x= 12+10= 22
----------------------------------------------
Check:
12*20= 240
22*10= 220
_____------
Total: 460 We have the correct answer